An object is 39 cm front of a converging lens with a focal length of 28 cm. What is the magnification of the lens


1
Expert's answer
2021-05-05T14:35:37-0400

Solution. Using lens formula


"\frac{1}{u}+\frac{1}{v}=\frac{1}{F}"

where u is object distance; v is image distance; F is focal length.



According to the condition of the problem u=0.39m, F=0.28m. Therefore


"\frac{1}{v}=\frac{1}{F}-\frac{1}{u}""v=\frac{Fu}{u-F}=\frac{0.39\times0.28}{0.39-0.28}\approx0.993m"

Using the figure, the magnification is


"M=\frac{h'}{h}=\frac{v}{u}=\frac{0.993}{0.39}\approx2.54"

Answer. M=2.54